OpenStudy (anonymous):

line AB is represented by the equation y=3x-1 line CDis parallel to line AB and passes through the points 5,1 what is the equation of line CDin slope intercept form?

OpenStudy (anonymous):

we know from the general form y=mx+b that m is our slope. In your given line, 3 is the slope. Therefore use the point-slope form (y-1)=3*(x-5) to get your slope-intercept form... y-1=3x-15 y=3x-14